# TAVP: Task-Adaptive Visual Prompt for Cross-domain Few-shot Segmentation

### v1

@@ III-A Problem Definition
In the field of cross-domain few-shot semantic segmentation (CD-FSS), we distinguish between a source domain $(X_{s},Y_{s})$. A target domain $(X_{t},Y_{t})$, with differing input distributions and non-overlapping label spaces, meaning $X_{s}\neq X_{t}andY_{s}\cap Y_{t}=\varnothing$. Here, `$X^{\prime}$ signifies the input distribution and `$Y^{\prime}$ denotes the label space. Our methodology involves training and evaluating our model episodically within a meta-learning framework as outlined in [4]. Training episodes consist of a support set and a query set. The support set $S$ = ${(I_{i}^{s},M_{i}^{s})}$ from $i=1$ to $K$ includes $K$ pairs of images and their respective binary masks, with $I$ representing the $i_{t}h$ support image and $M_{i}^{s}$ the matching binary mask. The query set, defined as $Q$ = ${(I_{i}^{q},M_{i}^{q})}$ from $i=1$ to $K$, operates similarly. The model is fed with the support set $S$ and a query set $I_{q}$ from a specific class $c$, upon which it is tasked with predicting the binary mask $M_{q}$ of the query. For testing, the model’s effectiveness is gauged by providing it with a new support set and corresponding queries from the target domain.
@@
Figure 3: The overall architecture of proposed TAVP network. First, multi-level features of support and query images are obtained from SAM encoder, then combined with original pre-trained weights on SA-1B dataset[10]. CDTAP is employed for fine-tuning and meta transform. Simultaneously, dense embedding:$F_{prompt}$ and image embeddings were acquired as the input that fed into decoder. At last, the mask decoder predicts the query image. The $L_{dem}$ loss is used for learnable prompts supervision and fine-tuning, and the $L_{seg}$ is used for supervising auto-prompts generation.
III-B Multi-level Features Fusion
@@
High-level Global Feature Representation. We propose an advanced approach to enhance the mask resolution in SAM by incorporating efficient token learning. Rather than utilizing the coarse masks generated by SAM directly, our method involves a High-level token alongside a novel mask prediction layer to produce higher-quality masks. In this method, we maintain the original mask decoder of SAM but augment it with a newly defined learnable High-level token with the size of $1\times 256$. This token is combined with the existing output tokens from SAM ($4\times 256$) and prompt tokens with the size of $N_{prompt}\times 256$), serving as the augmented input for the SAM mask decoder. Like the original output token function, the High-level token engages in self-attention with the other tokens and partakes in token-to-image and image-to-token attention processes within each attention layer for feature refinement. The High-level token utilizes a shared point-wise MLP across decoder layers. After two decoder layers, it comprehensively understands global image semantics and conceals mask information from other output tokens. A novel three-layer MLP is then employed to derive dynamic convolutional kernels from the enriched High-level Token, executing a spatial point-wise operation with the amalgamated High-level feature to generate superior-quality masks.
@@
Our approach trains only the High-level Token and its associated three-layer MLPs to correct inaccuracies in the mask produced by SAM without directly fine-tuning SAM or using a post-refinement network. This method stands in contrast to traditional approaches in high-quality segmentation models. Our extensive testing highlights two primary benefits of this efficient token learning technique: 1) It substantially elevates the mask quality of SAM with only a minimal increase in parameters, thus optimizing the training process in terms of time and data efficiency; and 2) Adaptive token and MLP components prevent overfitting, preserving SAM’s zero-shot segmentation performance on new images without knowledge loss.
@@
Global and Local Feature Fusion. Accurate segmentation requires input features with global semantic context and precise local boundaries. To enhance mask quality further, we augment the mask decoder features of SAM with both advanced object context and refined edge information. Rather than directly utilizing the mask decoder feature of SAM, we construct new multi-level features by extracting and integrating features from various stages of the SAM model. We first extract detailed low-level edge information from the initial layer’s local feature of SAM’s ViT encoder with a spatial dimension of 64x64. This feature is obtained from the first global attention block within the ViT encoder, specifically the 6th block of 24 blocks in case of a SAM based on ViT-Large. Then, the last layer’s high-level global feature from SAM’s ViT encoder, sized at $64\times 64$, provides a comprehensive global image context. Finally, the mask feature within SAM’s mask decoder, sized at $256\times 256$, is shared by the output tokens and possesses strong mask shape information. As depicted in Figure 3, we initially upsample the early and final layer encoder features to a spatial size of $256\times 256$ via transposed convolution to generate the input high-level features. Following this, we combine these three types of features through element-wise summation after straightforward convolutional processing. This approach of fusing global and local features is straightforward yet effective, producing segmentation results that preserve detail with minimal memory and computational costs. In the experimental section, we further conduct a detailed ablation study to assess the impact of each feature source.
III-C Task-adaptive Transferring (TAT) Class Domain Task-Adaptive Auto-Prompt (CDTAP)
@@
We improve the model’s generalization by performing class-domain prototype information disentanglement and using prior-guided automatic prompts for complete task-adaptive automatic prompts generation. The learnable prompt embedding increases the robustness of SAM and our model.
@@
Figure 4: (a)The previous class-wise few-shot methods. (b)Our two-way matching meta-learning module.
@@
Class Domain Prototype Information Disentanglement. Previous meta-learning methods only have generalizations for new categories, but the performance is poor when faced with both cross-domain and few-shot tasks. Therefore, we propose a new prototype-based class-domain information disentanglement module. To better explore the correlation between different class and domain features for distinct components, class-domain-common and class-domain-specific, we add a branch for the foreground and background prototype calculations. Using pixel-level prototype calculations can make fuller use of feature representation, which is helpful in few-shot learning. Previous methods only rely on the support prototype set and anchor layer to calculate the transformation matrix. Due to intra-class variance, the support prototype cannot represent all the information in the category. Therefore, we propose to enhance the set of supported prototypes by querying the prototypes. We specifically focus on dual prototype enhancement and cross-domain feature transformation. We leverage cycle consistency between support and query functions to obtain query foreground and background prototypes. Based on these enhanced prototypes that can represent categories and their surroundings, learnable domain-agnostic modules can be used to compute efficient transformation matrices. The transformation matrix is then applied to the query features for cross-domain feature transformation. Representational archetypes are important for our cross-domain transformation. To this end, we construct a pixel-level fine-grained self-cycling supervision that reasons the query foreground and background to support enhancement. We perform forward matching to obtain the query features with the highest similarity to the supporting prospects. We then use these identified forward-matching query features to backward retarget the corresponding supporting features. If the supporting features found by reverse matching fall within the true supporting foreground mask, the identified query features are averaged and used to derive the foreground prototype. An enhanced background prototype is obtained through the same process. Assuming that $W$ represents the weight matrix of original features in the Anchor layer and $P$ represents the prototype of the foreground representation. Specifically, we use equation 1 to get P from W and A. The difference between our algorithm and HQ-SAM [36] for Multi-level features lies in the design of CDTAP module, which transfers the original multi-scale based on feature computation to multi-scale based on prototype computation.
[…]
Where W is a learnable weight matrix, P is the computed prototypes and A is a representation matrix calculated from the distance between the center and other features. The prototype of foreground and background can be calculated by Equation  2
$P_{f,b}=\begin{bmatrix}\frac{P_{f}}{||P_{f}||},&\frac{P_{b}}{||P_{b}||}\end{bmatrix}$ (2)
$i^{s->q}=argmax(sim(P^{s}_{f}\odot P_{m}^{s}),P_{f}^{q})$ (3)
$j^{q->s}=argmax(sim(P^{q,f}_{i^{s->q}},P^{s,f}_{j^{q->s}}),$ (4)
@@
where $i$ and $j$ are rows and columns of 2D spatial positions of the feature map, Equation 3 and Equation 4 are the cycling check process. $P^{s}_{f}$ is the prototype representation of an image and $P_{m}^{s}$ is the prototype representation of its mask, $\odot$ represents the multiplication between vectors. $P^{q,f}_{i^{s->q}}$ is the feature prototype of the query from support to query matching. The corresponding interpretation can be deduced for $P^{s,f}_{j^{q->s}}$. Given the based equations, the prototype representation of the query can be selected.
@@
After the experiments, we prefer that not only the CDFSS task rely on better guidance and representation of foreground and essential information for the input sample, but also the background plays a critical role in more accurate predictions. For example, the more random the background setting is, or the greater the difference between the data in the source domain, the better the model performs.
III-D Light-Weight Fine-tune Framework
Besides, we adopt a random heterogenization sampling strategy to distinguish different cross-domain tasks. In this strategy, a threshold value is created to supervise the sampling quality. One of the limitations of SAM is that it is time-consuming and inefficient, which is a common issue in large model fine-tuning. In this work, we propose a lighted-weight fine-tuning framework, transferring SAM to cross-domain few-shot segmentation only by re-training a few layers in CNN-based models. First, the target domain samples are fed into the class domain task-specific branch for class-agnostic feature extraction. These highly structured class-agnostic feature embeddings, combined with other feature embeddings from the base domain, are fed into the decoder. A weighted supervision loss is proposed to fine-tune the decoder to predict masks for target domain samples. Lseg represents the segmentation loss function, composed of the Cross-Entropy loss function [37], and a Dice loss function [38], as defined in
@@ III-D Light-Weight Fine-tune Framework
where $\lambda$ is an adjustable parameter for supervision. Simultaneously, samples from the target domain are fed into a CNN-based model to generate dense embeddings as auto prompts. The dense embedding is obtained from the layer of CNN-based backbones as a weight matrix aligned with a feature map. Then, the dense embedding is multiplied with a combined multi-level feature map and is fed into a decoder, achieving guided decoding for target domain samples. Given an input $x$, it is fed into a CNN-based encoder. After down-sampling, a simple decoder follows for up-sampling to generate dense embedding, aligning with the feature map. The $L_{dem}$ loss function is adopted to supervise dense embedding, as shown in
@@
where $Z_{x}$ represents the dense embedding of input $x$, and $M_{x}$ is the mask of input $x$.
Overall, the end-to-end training framework is supervised by the following loss function
@@
where Lseg represents segmentation branch supervision and Ldem, is the loss function for dense embedding generating.
@@ IV-A1 Data Sets
We divide each image into six sections to increase the number of testing images and reduce their sizes. As the object categories in this data set have irregular shapes, cutting the images minimally impacts their segmentation. We further filter out images with only one class and those belonging to the ‘unknown’ category. This results in 5,666 images used to report the results, each with a resolution of $408\times 408$ pixels.
ISIC. The data set identified as ‘document number 1’, as described in references [40, 41], focuses on skin lesion imagery from cancer screenings, containing 2,596 images with a single lesion each. Ground-truth labels are provided solely for the training set. For consistent analysis, images are resized to a standard $512\times 512$ pixels from the original $1022\times 767$ pixels.
ChestX. As discussed in [42, 43], the Chest X-ray data set is tailored for Tuberculosis detection. It comprises a total of 566 X-ray images, each having an original resolution of $4020\times 4892$ pixels. These images are sourced from a data set of 58 cases with a Tuberculosis manifestation and 80 cases with normal conditions. Given the large size of the original images, a common practice is to reduce them to a more manageable $1024\times 1024$ pixels for further analysis and processing.
FS1000. FSS-1000 [44] is a natural image dataset for few-shot segmentation, consisting of 1,000 object classes, each with 10 samples. We used the official split for semantic segmentation in our experiment and reported the results on the official testing set, which contains 240 classes and 2,400 testing images.
@@
In this work, a heterogenization sampling strategy is adopted to reduce the coupling effect of training with a limited data set. In this strategy, a threshold is set to supervise the sampling quality. Specifically, we used a 5-fold validation strategy in the split data set once during the training process. The threshold is computed through the training process to select the samples for model training dynamically.
Besides, several data augmentation methods also used in original SAM baseline are utilized, including adjusting the attributes of images such as brightness, contrast, saturation, etc., randomly flipping along vertical and horizontal levels, and randomly affine transformation.
IV-A3 Models
The overall framework is based on High-quality Segment Anything[36]. Simultaneously, a task-adaptive auto-prompts branch, together with SAM, is proposed to complete the CDFSS.
IV-A4 Training Setting
For the training data set, only one image from the target domain including ChestX [2, 15], ISIC [9], FSS-1000 [22], and deepglobe [10] at a time are used respectively with augmentation. For training parameters, the baseline is frozen, which shown as the snowflake logo on Figure3. only the CDTAP module is training.
IV-A5 Implementation Details.
We have designed three backbones for the additional branch except the SAM framework. SAM encoder was adapted for high-level global semantic contexts and low-level local information computing. The full ViT is used for computing in global semantic contexts, and the local low-level features are extracted from the early layer. Besides, a CNN-based encoder, modified for computing prototypes, is used for class-domain agnostic feature extraction, and a small up-sampling method is adopted for dense embedding generation as auto-prompts.
During the training stage, the number of epochs is between 60 and 150. Since there is a large weight pre-trained SAM model, the epochs are smaller for end-to-end training to avoid overfitting. Our experiments can achieve ideal results if the running time is between 2 and 6 hours on NVIDIA A6000 GPU, depending on the number of epochs and validation data sets.
V Visualizations
Figure 6: Qualitative results of TAVP in 1-way 5-shot segmentation on CD-FSS. Support labels are overlaid in red. The ground truth and predictions of query images are in green and blue, respectively.
TABLE I: Comparison with previous FSS and CD-FSS methods under 1-way 1-shot and 5-shot settings on the CD-FSS benchmark.
@@ V Visualizations
ViT Based and Cross Domain Few Shot Segmentation Methods
@@
V-A Comparison with SOTA Methods
Extensive experiments are conducted to compare our method with the state-of-the-art methods. The results showed that we achieved better results on the Deepglobe data set than the latest SOTA performance. Besides, in the other three cross-domain data sets, we achieve better and more competitive and accurate results than the previous methods shown in Table I. Moreover, it is obvious that with a more robust model and flexible learning ability, the prediction is closer to the samples’ original semantics, especially pixel-level information, instead of relying on fixed ground truth, as shown in Figure 6.
V-B Ablation Study
Given the proposed method, we test the performance of models with different combination strategies. Overall, we divide the ablation experiment into the following parts based on different backbones, data augmentation strategy, fusion branches, and ablation study with SOTA. All ablation experiments are based on the pre-trained weight: ‘vit_h’ for better performance. Besides, the FS1000 data set has very little cross-domain difficulty and is not representative, so we did not use it in ablation experiments to test the effectiveness of our algorithms but only used it in comparative experiments in Table I.
Backbone and Data Augmentation Ablation. In the additional task-specific class-domain agnostic feature extraction and auto-prompts generation branch, we performed ablation experiments on the models’ performance on three data sets that are more difficult to cross-domain under 1-way 1-shot and 5-shot settings, and the details are shown in Table II. This study proves that ResNet has a stronger recognition ability for categories, and its effect is outstanding in our novel learnable prompt.
TABLE II: Ablation Study of different backbones under 1-way 1-shot and 5-shot settings.

### v2

@@ III-A Problem Definition
In the field of cross-domain few-shot semantic segmentation (CD-FSS), let $X_{s}$ and $X_{t}$ stand for the input distributions in the source and the target domains, respectively, and $Y_{s}$ and $Y_{t}$ denote the label spaces in the two domains, respectively. We distinguish between $\{X_{s},Y_{s}\}$ and $\{X_{t},Y_{t}\}$ with differing input distributions and non-overlapping label spaces, i.e., $X_{s}\neq X_{t}$ and $Y_{s}\cap Y_{t}=\varnothing$. Our methodology involves training and evaluating our model episodically within a meta-learning framework as outlined in [6]. Training episodes consist of a support set and a query set. The support set $S=\{(I_{i}^{s},M_{i}^{s})|i=1,2,\cdots,K\}$, where $I_{i}^{s}$ is the $i$-th support image and $M_{i}^{s}$ is the respective binary mask. The query set, defined by $Q=\{(I_{i}^{q},M_{i}^{q})|i=1,2,\cdots,K\}$, operates similarly. The model is fed with the support set $S$ and a query image set $I^{q}=\{I_{i}^{q}\}_{i=1}^{K}$, from a specific class $c$, upon which the binary mask set $M^{q}=\{M_{i}^{q}\}_{i=1}^{K}$ is predicted.
III-B Method Overview
While SAM can be generalized to more scenarios, even zero-shot situations, it still has some limitations. Firstly, the original SAM relies on interactive prompts for accurate segmentation in different situations, which can be time-consuming. The second challenge is how to transfer richer knowledge and key information from LVM methods while maintaining strong generalization ability. To address these two challenges, we propose an automated framework for segmentation that uses automatic prompts instead of user-interactive prompts. Additionally, we have designed an extra branch for class- and domain-agnostic feature extraction and task-adaptive prompt generation.
The overall framework for TAVP is shown in Figure 2. The inputs of images from the source domain with cut-mix are fed into the SAM encoder for basic feature extraction. Note that we propose a multi-level feature fusion for extensive representation. Meanwhile, one pair of support and query images from the target domain are fed into the CDTAP module for class domain-specific and agnostic feature extraction. At the same time, this module generates learnable prompts as dense embedding input to the decoder. Then, the combined multi-level and dense prompts are fed into the SAM decoder for prediction.
@@ III-B Method Overview
Figure 2: The overall architecture of proposed TAVP network. First, the images from source domain with cut-mix are passed through the SAM encoder to obtain multi-level features, which are combined with original pre-trained weights on SA-1B dataset [12], and followed by a batch normalization layer to get the class-agnostic features. Additionally, CDTAP is employed for fine-tuning and meta-transformation. Simultaneously, dense embedding: $F_{prompt}$ and image embeddings are acquired as the input of the decoder. At last, the mask decoder predicts the query image. The $L_{dem}$ loss is used for learnable prompt supervision and fine-tuning, and $L_{seg}$ is used for supervising auto-prompts generation.
III-C Multi-level Features Fusion
@@
High-level Global Feature Representation. We propose an advanced approach to enhance the mask resolution in SAM by incorporating efficient token learning. Rather than utilizing the coarse masks generated by SAM directly, our method involves a High-level token alongside a novel mask prediction layer to produce higher-quality masks. In this method, we maintain the original mask decoder of SAM but augment it with a newly defined learnable High-level token with the size of $1\times 256$. This token is combined with the existing output tokens with the size of ($4\times 256$) from SAM and prompt tokens with the size of $N_{prompt}\times 256$), serving as the augmented input for the SAM mask decoder. Like the original output token function, the High-level token engages in self-attention with the other tokens participating in token-to-image and image-to-token attention processes within each attention layer for feature refinement. The High-level token utilizes a shared point-wise MLP across decoder layers. After two decoder layers, it comprehensively understands global image semantics and conceals mask information from other output tokens. A novel three-layer MLP is then employed to derive dynamic convolutional kernels from the enriched High-level Token, executing a spatial point-wise operation with the amalgamated High-level feature to generate superior-quality masks.
@@
Our approach trains only the High-level token and its associated three-layer MLPs to correct inaccuracies in the mask produced by SAM without directly fine-tuning SAM or using a post-refinement network. This method stands in contrast to traditional approaches in high-quality segmentation models. Our extensive testing highlights two primary benefits of this efficient token-learning technique. First, it substantially elevates the mask quality of SAM with only a minimal increase in parameters, thus optimizing the training process in terms of time and data efficiency. Second, adaptive token and MLP components prevent overfitting, preserving SAM’s zero-shot segmentation performance on new images without knowledge loss.
@@
Global and Local Feature Fusion. Accurate segmentation requires input features with global semantic context and precise local boundaries. To enhance mask quality further, we augment the mask decoder features of SAM with both advanced object context and refined edge information. Rather than directly utilizing the mask decoder feature of SAM, we construct new multi-level features by extracting and integrating features from various stages of the SAM model. We first extract detailed low-level edge information from the initial layer’s local feature of SAM’s ViT encoder with a spatial dimension of $64\times 64$. This feature is obtained from the first global attention block within the ViT encoder, specifically the 6th block of 24 blocks in the case of a SAM based on ViT-Large. Then, the last layer’s high-level global feature from SAM’s ViT encoder, sized at $64\times 64$, provides a comprehensive global image context. Finally, the mask feature within SAM’s mask decoder, sized at $256\times 256$, is shared by the output tokens and possesses strong shape information of the masks. As depicted in Figure 2, we initially upsample the early and final layer encoder features to a spatial size of $256\times 256$ via transposed convolution to generate the input high-level features. Following this, we combine these three types of features through element-wise summation after straightforward convolutional processing. This approach of fusing global and local features is straightforward yet effective, producing segmentation results that preserve detail with minimal memory and computational costs. In the experimental section, we further conduct a detailed ablation study to assess the impact of each feature source.
III-D Class Domain Task Adaptive Auto Prompt (CDTAP)
@@
We improve the model’s generalization by disentangling class-domain prototype information and using prior-guided prompts for fully automatic, task-adaptive prompt generation. The learnable prompt embedding increases the robustness of SAM and our model.
@@
Figure 3: Existing class-wise few-shot methods and our two-way matching meta-learning module.
Figure 4: Details of MT and PG.
Class Domain Prototype Information Disentanglement. Previous meta-learning methods only have generalizations for new categories, but the performance degrades when handling both cross-domain and few-shot tasks. To address this, we propose a new prototype-based class-domain information disentanglement module. This module aims to better explore the correlation between class and domain features by separating them into class-domain-common and class-domain-specific components. The foundation segmentation model [12] is used for base knowledge regulation and a branch for the foreground and background prototype calculations is added. Pixel-level prototype calculations fully utilize feature representations, which is beneficial for few-shot learning.
The pre-trained foundation segmentation model contains a large number of base class knowledge. We extract low-level features for wider expression and high-level features for semantic expression. Then, a batch normalize layer is used for regularization to get class-agnostic knowledge.
@@
Previous methods only rely on the support prototype set and anchor layer to calculate the transformation matrix. Due to intra-class variance, the support prototype cannot represent all the information in the category. Therefore, we propose to enhance the set of supported prototypes by querying the prototypes. We specifically focus on dual prototype enhancement and cross-domain feature transformation. We leverage cycle consistency between support and query functions to obtain query foreground and background prototypes. Based on these enhanced prototypes that can represent categories and their surroundings, learnable domain-agnostic modules can be used to compute efficient transformation matrices. The transformation matrix is then applied to the query features for cross-domain feature transformation. Representational archetypes are important for our cross-domain transformation. To this end, we construct a pixel-level fine-grained self-cycling supervision that reasons the query foreground and background to support enhancement. We perform forward matching to obtain the query features with the highest similarity to the supporting prospects. We then use these identified forward-matching query features to backward retarget the corresponding supporting features. If the supporting features found by reverse matching fall within the true supporting foreground mask, the identified query features are averaged and used to derive the foreground prototype. An enhanced background prototype is obtained through the same process. Let $W$ represent the weight matrix of original features in the Anchor layer and $P$ represent the prototype of the foreground representation. Specifically, we use $WP=A$ to obtain $P$ from $W$ and $A$ [6], where $W$ is a learnable weight matrix, $P$ is the computed prototypes and $A$ is a representation matrix calculated from the distance between the center and other features. The difference between our algorithm and HQ-SAM [44] for Multi-level features lies in the design of the CDTAP module, which transfers the original multi-scale based on feature computation to multi-scale based on prototype computation. The prototype of foreground and background can be calculated by Equation  1
$P_{f,b}=\begin{bmatrix}\frac{P_{f}}{||P_{f}||},&\frac{P_{b}}{||P_{b}||}\end{bmatrix},$ (1)
$i^{s->q}=\text{argmax}(\text{sim}(P^{s}_{f}\odot P_{m}^{s}),P_{f}^{q}),$ (2)
$j^{q->s}=\text{argmax}(\text{sim}(P^{q,f}_{i^{s>q}},P^{s,f}_{j^{q->s}})),$ (3)
@@
where $i$ and $j$ are rows and columns of 2D spatial positions of the feature map. Equation 2 and Equation 3 are the cycling check process, where $P^{s}_{f}$ is the prototype representation of an image, $P_{m}^{s}$ is the prototype representation of its mask, $\odot$ represents the multiplication between vectors, and $P^{q,f}_{i^{s->q}}$ is the feature prototype of the query from support to query matching. The corresponding interpretation can be deduced for $P^{s,f}_{j^{q->s}}$. Given the based equations, the prototype representation of the query can be selected.
@@
III-E Light-Weight Fine-tune Framework
Besides, we adopt a random heterogenization sampling strategy to distinguish different cross-domain tasks. In this approach, a threshold value is set to monitor the quality of the sampling process. One of the limitations of SAM is that it is time-consuming and inefficient, which is a common issue in large model fine-tuning. In this work, we propose a lightweight fine-tuning framework, transferring SAM to cross-domain few-shot segmentation only by re-training a few layers in CNN-based models. First, the target domain samples are fed into the class domain task-specific branch for class-agnostic feature extraction. These highly structured, class-agnostic feature embeddings, together with other feature embeddings from the base domain, are fed into the decoder. A weighted supervision loss is proposed to fine-tune the decoder to predict masks for target domain samples. Lseg represents the segmentation loss function, composed of the Cross-Entropy loss function [45], and a Dice loss function [46], as defined in
@@ III-E Light-Weight Fine-tune Framework
where $\lambda$ is an adjustable parameter for supervision. Simultaneously, samples from the target domain are fed into a CNN-based model to generate dense embeddings as auto prompts. The dense embedding is obtained from the layer of CNN-based backbones as a weight matrix aligned with a feature map. Then, the dense embedding is multiplied with a combined multi-level feature map and is fed into a decoder, achieving guided decoding for target domain samples. Given an input $x$, it is fed into a CNN-based encoder. After down-sampling, a simple decoder follows for up-sampling to generate dense embedding, aligning with the feature map. The $L_{dem}$ loss function is adopted to supervise dense embedding:
@@
where $Z_{x}$ represents the dense embedding of input $x$, and $M_{x}$ is the mask of input $x$. Overall, the end-to-end training framework is supervised by the following loss function
@@
where Lseg represents segmentation branch supervision and Ldem is the loss function for dense embedding generating.
@@ IV-A Experimental Settings
Follow the standard approach to previous work [6], We divide each image into six sections to increase the number of testing images and reduce their sizes. Since the object categories in this dataset have irregular shapes, cutting the images has minimal impact on their segmentation. We further filter out images with only one class and those belonging to the ‘unknown’ category. This results in 5,666 images used to report the results, each with a resolution of $408\times 408$ pixels.
ISIC. The dataset identified as ‘document number 1’, as described in [48, 49], focuses on skin lesion imagery from cancer screenings, containing 2,596 images with a single lesion. Ground-truth labels are provided solely for the training set. Follow the standard approach to previous work [6], for consistent analysis, images are resized to a standard $512\times 512$ pixels from the original $1022\times 767$ pixels.
ChestX. As discussed in [50, 51], the Chest X-ray dataset is tailored for Tuberculosis detection. It comprises a total of 566 X-ray images, each with an original resolution of $4020\times 4892$ pixels. These images are sourced from a dataset of 58 cases with a Tuberculosis manifestation and 80 cases with normal conditions. Given the large size of the original images, a common practice is to reduce them to a more manageable $1024\times 1024$ pixels for further analysis and processing.
FS1000. FSS-1000 [52] is a natural image dataset for few-shot segmentation, consisting of 1,000 object classes, each with 10 samples. We use the official split for semantic segmentation in our experiment and report the results on the official testing set, which contains 240 classes and 2,400 testing images.
@@
In this work, a cut-mix method and a heterogenization sampling strategy are adopted to reduce the coupling effect of training with a limited dataset. First, images from the target domain are randomly divided into patches of different sizes, and then the image from the original domain is used as the background to create a new input image. In the experiment, each newly synthesized image contains 5 patches from the original domain. In the second strategy, a threshold is set to control the quality of the sampling. Specifically, we applied a 5-fold validation strategy on the split dataset during the training process. The threshold is computed dynamically during training to select samples for model training. The size of each patch is 4892 pixels.
Besides, several data augmentation methods are also used in the original SAM baseline, including adjusting the attributes of images such as brightness, contrast, saturation, etc., randomly flipping along vertical and horizontal levels, and random affine transformation.
After the experiments, we emphasize the importance of ensuring that the CD-FSS task relies on improved guidance and representation of foreground and essential information for the input sample, while highlighting that the background plays a critical role in achieving more accurate predictions. This emphasizes the efficiency of our method as shown in Figure 3. For example, the more random the background setting is, or the greater the difference between the data in the source domain is, the better the model performs.
TABLE I: Comparison with previous FSS and CD-FSS methods under 1-way 1 shot and 5-shot settings on the CD-FSS benchmark.
@@
ViT Based Methods and Cross Domain Few Shot Segmentation Methods
@@
IV-A3 Implementation Details.
We have designed three backbones for the additional branch except the SAM framework. SAM encoder is adapted for high-level global semantic contexts and low-level local information extraction. The full ViT is used for global semantic contexts, and the local low-level features are extracted from the early layer. Besides, a CNN-based encoder, modified for computing prototypes, is used for class-domain agnostic feature extraction, and a small up-sampling method is adopted for dense embedding generation as auto-prompts.
There are two parts of input here. The first part uses enhanced data, a cutmix library (a total of 1,000 images) is generated to increase feature diversity. We randomly cropped the images in the target domain into different patches and pasted them on the original domain images. Then the augmented datasets are fed into SAM encoder to extract image embedding for late decoding. For the another part, only one pair of support and query images from the target domain (including ChestX [50], ISIC [48], FSS-1000 [52] and deepglobe [47]) are fed into the CDTAP module. For the training parameters, the baseline is frozen, indicated by the snowflake icon in Figure 2, and only the CDTAP module is trained.
In the training stage, the number of epochs is set to be between 60 and 150. Our experiments can achieve ideal results if the running time is between 2 and 6 hours on NVIDIA A6000 GPU, depending on the number of epochs and validation datasets.
Figure 5: Qualitative results of TAVP in 1-way 5-shot segmentation on CD-FSS. Support labels are overlaid in red. The ground truth and predictions of query images are highlighted, respectively.
IV-B Comparison with SOTA Methods
Extensive experiments are conducted to compare our method with the state-of-the-art methods. The results show that we achieve better results on the Deepglobe dataset than the latest SOTA performance. Besides, in the other three cross-domain datasets, we achieve better and more competitive and accurate results than the previous methods as shown in Table I. Moreover, it is obvious that with a more robust model and flexible learning ability, the prediction is closer to the samples’ original semantics, especially pixel-level information, instead of relying on fixed ground truth, as shown in Figure 5.
IV-C Ablation Study
Given the proposed method, we test the performance of models with different combination strategies. Overall, we divide the ablation experiment into the following parts based on different backbones, data augmentation strategy, fusion branches, and ablation study with SOTA. All ablation experiments are based on the pre-trained weight: ‘vit_h’ for better performance. Besides, the FS1000 dataset presents minimal cross-domain difficulty and is not representative, so we do not use it in ablation experiments to test the effectiveness of our algorithms but only use it in comparative experiments in Table I.
TABLE II: Ablation Study of different settings under 1-way 1-shot and 5-shot settings.